OpenStudy (96mertcank96):

The quotient of (9.2 x 10^6)and (2.3 x 10^2) expressed in scientific notation is?

OpenStudy (anonymous):

\[\frac{ 9.2*10^{6} }{ 2.3*10^{2} }=\frac{ 9.2*10^{4} }{ 2.3 }=4*10^{4}=40000\]

OpenStudy (anonymous):

\[4*10^{4}\] Is the scientific notation.